(1) Dr Mukonda’s performance of his 170 Biostatistics class of students is given in an incomplete distribution below.


Variable 0-10 10-20 20-30 30-40 40-50 50-60 60-70

Frequency f1 21 32 40 21 26 f7


(a) If the median is 35. After finding the missing frequencies, find the standard deviation and present this information

on a pie chart.


(b) Statistically, explain if these results are normally distributed.


1
Expert's answer
2021-04-25T14:34:24-0400

Class interval Frequency Cumulative Frequency 

 0−10 10 10 

 10−20 20 30 

 20−30 "f_1" 30+"f _1"

 30−40 40 70+"f_1"

 40−50 "f_2"  70+"f _1+f_ 2"

​ 50−60 25 95+"f_1+f_2"

 60−70 15 110+"f _1+f_2"

  

(a) Here, Median=35.


∴ median class is 30−40,N=170,then "\dfrac{N}{2}=\dfrac{170}{2}=85"


"L=30,f=40,F=30+f_ 1,h=10"


⇒ Now,  Median"=L+\dfrac{\dfrac{N}{2}-F}{f}\times h"


"35=30+\dfrac{85-(30+f_1)}{40}\times 10"


 "\Rightarrow 35\times 4=120+(55-f_1)"


"\Rightarrow f_1=35"


 "Also f_2=170-110-f_1=60-35=25"


Hence, the missing frequencies are,

Class 20−30⇒35

Class 40−50⇒25


The table for the above data distribution is-




Mean "\mu=\dfrac{\sum XF}{\sum F}=\dfrac{6100}{170}=35.88"



Standard deviation "s=\sqrt{\dfrac{\sum (x-\mu)^2}{N}}=\sqrt{\dfrac{3016.59}{170}}=\sqrt{17.25}=4.212"


The pie chart is-




(b) The given data is not normally distributed since, There is no decrement in the frequencies after the middle frequency.